OpenStudy (anonymous):

Find the sum: -21 + (-13) + (-9).

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (solomonzelman):

A\[a+(-b)=a-b\]

OpenStudy (anonymous):

???

OpenStudy (solomonzelman):

for any numbers,\[a+(-b)=a-b\]apply this here.

OpenStudy (skullpatrol):

Since you are adding all negative numbers, add them as if they were all positive and then prefix their common sign.

OpenStudy (anonymous):

A. -43 B. -25 C. 1 D. 4

OpenStudy (solomonzelman):

\[-a+(-b)=-a-b\] \[-21 + (-13)=?\]

OpenStudy (anonymous):

−34

OpenStudy (skullpatrol):

$$\huge -21 + (-13) + (-9)= -[21 + 13 + 9]$$

OpenStudy (anonymous):

wdf i am confused

OpenStudy (solomonzelman):

I can't msg you, @skullpatrol, so can you tell me how you type this big numbers? I'l give you a medal!

OpenStudy (skullpatrol):

$$\huge \text{if a and b are both negative, then}$$ $$\huge a+b=-(|a|+|b|)$$

OpenStudy (skullpatrol):

$$\huge\text{Since you are adding all negative numbers,}$$ $$\huge\text{add them as if they were all positive}$$ $$\huge\text{ and then prefix their common sign.}$$

OpenStudy (solomonzelman):

I still don't get it I keep getting \[$$\huge$SEE\]

OpenStudy (solomonzelman):

@skullpatrol

OpenStudy (solomonzelman):

$$\huge$$See?